OpenStudy (anonymous):

Find the slope of the line that passes through (3, -4) and (-1, 2).

OpenStudy (anonymous):

slope is \[\frac{y2 - y1}{x2 - x1}\]

OpenStudy (anonymous):

so you do \[\frac{2 - (-4)}{-1 - 3}\]

OpenStudy (anonymous):

Simplify and you get \[- \frac{6}{4}\] Which simplifies down to \[- \frac{3}{2}\]

OpenStudy (anonymous):

slope is -3/2...